OpenStudy (anonymous):

a store that sells handcrafted items gets $3.00 + 40% for each item sold. The rest goes to the craftperson. All items sold Cost $5.00 or more. if p represents the sale price of one item what would represent what the amount of money craft person makes?

OpenStudy (anonymous):

\[p-\left(3+\frac{4}{10}p\right)=\frac{3 p}{5}-3 \]